# On Subjective Uncertainty Quantification and Calibration in Natural Language Generation

### v1

@@ Abstract
Applications of large language models often involve the generation of free-form responses, in which case uncertainty quantification becomes challenging. This is due to the need to identify task-specific uncertainties (e.g., about the semantics) which appears difficult to define in general cases. This work addresses these challenges from a perspective of Bayesian decision theory, starting from the assumption that our utility is characterized by a similarity measure that compares a generated response with a hypothetical true response. We discuss how this assumption enables principled quantification of the model’s subjective uncertainty and its calibration. We further derive a measure for epistemic uncertainty, based on a missing data perspective and its characterization as an excess risk. The proposed measures can be applied to black-box language models. We demonstrate the proposed methods on question answering and machine translation tasks, where they extract broadly meaningful uncertainty estimates from GPT and Gemini models and quantify their calibration.
@@ 1 Introduction
We are interested in uncertainty quantification for language models (LMs) in free-form natural language generation (NLG): given instruction $I$, the model generates a response $y^{\prime}$ based on its predictive distribution $p_{M}(y\mid I)$, where $y^{\prime}$ could be any natural language passage that fits the instruction. An example of this is question answering (QA): given a question from the user, the model may provide a brief answer, but it may also follow with supporting facts and explanations, which can vary in form and detail. The user can be satisfied by a wide variety of responses, irrespective of their style or (to some extent) the choice of supporting facts included.
@@
This paper is about the observation that the above challenges can be resolved in a more general setup from a perspective of Bayesian decision theory (Savage,, 1954). We first assume that we can compute a similarity-type measure $S(y^{\prime},y;I)$ that characterizes our utility function when the model generates $y^{\prime}$ in respond to $I$ and $y$ is a (true or hypothetical) correct response. This generalizes previous works based on semantic equivalence, which corresponds to defining $S$ using entailment. It can also be applied to traditional structured prediction tasks where a choice of $S$ is readily available (e.g., the chrF score for machine translation, Popović,, 2015), as well as tasks where evaluation can be adequately implemented using frontier language models that computes $S$ based on few-shot demonstrations or detailed instructions.
@@
To quantify the subjective uncertainty defined by the LM, we assume the generation $y$ is always chosen to maximize the expected utility under the LM’s predictive distribution $p_{M}$2.1). This is a common assumption in NLG (Bickel and Doksum,, 1997; Bertsch et al.,, 2023). It can be supported by a common belief that high-capacity LMs may approximate Bayesian inference (e.g., Xie et al.,, 2021; Akyürek et al.,, 2022; Hahn and Goyal,, 2023; Ye et al.,, 2024), and is also more broadly applicable whenever we view $p_{M}$ as our best available model for $y\mid I$. Subjective uncertainty is then naturally characterized by the Bayes risk, or equivalently the maximum achievable expected utility, where the action space is defined by the candidate generations available. This simple observation allows us to understand that previous methods for “semantic uncertainty” can be adapted to a broader range of scenarios, regardless of whether semantic equivalence is relevant. It also provides a unique, principled approach for aggregating similarity measures among generations (§2.2).
@@
For subjective uncertainty measures to be useful, it will be ideal if the LM is calibrated. It has been unclear in previous works how calibration can be evaluated in free-form NLG tasks, given the distinction between relevant and irrelevant differences for a specific task and our apparent inability to formally characterise it. The decision-theoretic view provides a natural answer to this question: an LM is deemed as to possess a calibrated notion of uncertainty if the expected subjective utility of its actions matches the expectation of the actually incurred utility, defined w.r.t. responses from the true data distribution (§2.3). This observation allows us to quantify the calibration of LMs through reliability diagrams (Murphy and Winkler,, 1977) and a generalized version of expected calibration error (Naeini et al.,, 2015). In the discussion we also note a generally applicable definition for “task-specific semantics”: the equivalence class defined by the action-utility map, known as Savage acts.
@@
Bayesian modelling allows for the decomposition of predictive uncertainty into epistemic uncertainty and aleatoric uncertainty (Der Kiureghian and Ditlevsen,, 2009). The decomposition allows us to better understand and diagnose the model’s uncertainty (e.g., only the epistemic uncertainty should decrease as the training set grows); it can also guide data acquisition (Kendall and Gal,, 2017). The quantification of epistemic uncertainty in general NLG applications appears challenging, due to the black-box nature of LMs and the free-form generations involved. §2.4 demonstrates that the decision-theoretic view enables us to quantify epistemic uncertainty for in-context learning (ICL, Brown et al.,, 2020) applications; key additional ingredients include a missing data perspective to Bayesian modelling (Fong et al.,, 2024) and a connection between epistemic uncertainty and a notion of excess risk (Xu and Raginsky,, 2022). The resulted measure is connected to and generalizes various previous works on epistemic uncertainty quantification for black-box models, which provides independent justification for its use.
@@
The rest of this paper is structured as follows: §2 presents the proposed methodology and discusses connections to previous work; §3 provides experimental illustrations; and §4 provides concluding remarks.
2 Quantifying Subjective Uncertainty and Calibration
@@ 2.1 A Utilitarian Setup
The basic setup of this paper is as follows:
@@
  • our utility can be measured through the similarity measure $S(y^{\prime},y;I)\in\mathbb{R}$, which can be cheaply evaluated;
  • we have access to an LM $p_{M}$, and, in the absence of further evidence, consider $p_{M}$ sufficiently trustworthy so that the ideal generation (i.e., action) $y^{\prime}$ given $I$ should maximize the expected utility,
@@
    where $\mathcal{Y}^{\prime}_{I}\subset\mathcal{Y}$ denotes the space of candidate generations.
In the above, $p_{M}$ represents a subjective belief about the true data distribution. $\mathcal{Y}_{I}^{\prime}\subset\mathcal{Y}$ can be determined based on computational constraints, e.g., as a separate set of samples from $p_{M}(\cdot\mid I)$.111 When $\mathcal{Y}_{I}^{\prime}$ is stochastic, expected utility/risk quantities such as (1) should also be averaged over the randomness in $\mathcal{Y}_{I}^{\prime}$. For brevity, we omit such (outmost) expectations throughout the paper. The utility $S$ determines a risk function $r(y^{\prime},y;I)=-S(y^{\prime},y;I)$, which can be scaled and shifted as desired.
As discussed in §1, this setup is very general. This decision-theoretic perspective is taken in the framework of minimum Bayes risk (MBR) decoding (Bickel and Doksum,, 1997), which advocates for generating $y$ based precisely on (1). MBR recovers many modern methods for natural language generation (Bertsch et al.,, 2023), and these methods can be viewed as having adopted the same view implicitly. Note that adopting the decision-theoretic setup does not compel us to assume that $p_{M}$ is a strictly Bayesian model (e.g., in the sense of Xie et al.,, 2021): although that will provide a good justification, it is also possible that a non-Bayesian LM may better approximate our belief than any available Bayesian model does.
2.2 Task-Specific Measure of Subjective Uncertainty
@@
$R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I}):=\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}r(y^{\prime},y;I).$ (2)
For example, if $\mathcal{Y}=\mathcal{Y}^{\prime}_{I}\subset\mathbb{R}$ and $r(y,y^{\prime};I)=(y-y^{\prime})^{2}$, $R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ will be equivalent to the variance of the predictive distribution; if with an abuse of notation we redefine $y^{\prime}$ as a probability density function and $r(y^{\prime},y;I)\leftarrow\log y^{\prime}(y)$, $R_{B,\mathcal{Y}^{\prime}_{I}}(I)=H[p_{M}(\cdot\mid I)]$ will be the entropy. When the action space $\mathcal{Y}^{\prime}_{I}\subsetneq\mathcal{Y}$ is a strict subset, (2) will be connected to measures of usable information as discussed in Xu et al., (2019).
It is thus natural to also adopt (2) as a measure of subjective uncertainty for NLG tasks. To implement (2), we adopt Monte Carlo approximation for the expectation.
A common generation strategy is to follow the Gibbs predictor, which corresponds to (1) with $\mathcal{Y}^{\prime}_{I}:=\{y_{I}\}$ and $y_{I}\sim p_{M}(\cdot\mid I)$ being a single sample. In such cases, the expected uncertainty $\mathbb{E}_{\mathcal{Y}^{\prime}_{I}}R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ is connected to the degree statistic in Lin et al., (2024, Eq. 8); the difference is that we use a similarity measure that is assumed to define our utility. While Lin et al., (2024) explored 9 different uncertainty measures for the Gibbs predictor, the decision-theoretic perspective leads to one unique method without any underspecified design choices, justifies its use beyond the scope of semantic equivalence, and highlights that different choices of predictors ($\mathcal{Y}_{I}^{\prime}$) should be matched with different uncertainty measures. However, being a measure of subjective uncertainty, it is not guaranteed to always improve on downstream tasks compared with heuristic alternatives, especially if the LM is less calibrated.
@@ 2.3 Evaluation of Task-Specific Calibration
Subjective uncertainty measures are most useful when the LM is calibrated. As discussed in Huang et al., (2024), many previous works studying uncertainty on NLG tasks did not distinguish between calibration and predictive performance in their evaluation, which leads to a relative lack of understanding about the calibration of LMs in such scenarios.
A typical definition of calibration is as follows (see e.g., Johnson et al.,, 2024): suppose the spaces $\mathcal{Y},\mathcal{I}$ are discrete, and denote the true data distribution as $p_{0}$; a model $p_{M}$ is considered calibrated w.r.t. some grouping function $G(I):\mathcal{I}\to\mathcal{G}$ if we have
@@
The grouping function $G$ may be defined based on the model’s prediction, or it may incorporate additional information such as input demographics. In the NLG setting, however, the criterion (3) can be trivially violated if, e.g., the LM always generates a longer response compared with $p_{0}$.
In light of the assumptions in §2.1, it is natural to relax (3) as follows:
$\mathbb{E}(p_{M}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$ $=\mathbb{E}(p_{0}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$
$\hskip 130.0002pt~{}~{}\forall g\in\mathcal{G},y^{\prime}\in\mathcal{Y}^{\prime}_{I},s\in\mathbb{R}.$ (3’)
@@
$s_{y,I}:\mathcal{Y}_{I}^{\prime}\to\mathbb{R},~{}y^{\prime}\mapsto S(y^{\prime},y;I)$ (4)
is injective (e.g., if $S$ is defined by a metric on $\mathcal{Y}$). (3) becomes weaker if $S(y^{\prime},y;\mathcal{I})$ only depends on $y$ and $y^{\prime}$ through a non-invertible function, e.g., a map from a natural language passage to its semantics; in such cases, it is clear that (3) remains a sufficiently strong criterion for task-specific calibration. More generally, we can always view the function $s_{y,I}(y^{\prime})$ as a representation of “task-specific semantics”.222 For the goals of providing the best generation from $\mathcal{Y}^{\prime}_{I}$ and quantifying its risk, it suffices to define the function $s_{y,I}$ on $\mathcal{Y}^{\prime}_{I}$. For purposes beyond the scope of the generation task, it may be helpful to extend its domain to a larger subset of $\mathcal{Y}$. In fact, the replacement of $y^{\prime}\in\mathcal{Y}^{\prime}_{I}$ with the more compact $s_{y^{\prime},I}(\cdot)$ has been advocated by Savage, (1954), and $s_{y,I}$ is referred to as a Savage act (see e.g., Marinacci,, 2015, §2.4).
A main reason we are interested in criteria such as (3) or (3) is that they guarantee the calibration of uncertainty measures. It follows from (3) that $\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(y^{\prime},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(y^{\prime},y;I)$ for all $y^{\prime}$, and thus
$\mathbb{E}_{p_{0}(I\mid G(I)=g)}R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(\hat{y}_{I},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(\hat{y}_{I},y;I),$ (5)
where $\hat{y}_{I}$ denotes the MBR generation (1). Plugging in $G(I)=R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$, we find that LMs are calibrated in the sense of (3) only if
$\text{gECE}(p_{M}):=\mathbb{E}_{s}|f_{M}(s)-s|=0,~{}~{}~{}\text{where }f_{M}(s):=\mathbb{E}_{I,y,\hat{y}_{I}}(r(\hat{y}_{I},y;I)\mid R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})=s)~{}\forall s\in\mathbb{R}.$ (6)
(6) provides a natural generalization of the expected calibration error (ECE, Naeini et al.,, 2015) in classification, which it recovers with $r(y^{\prime},y;I)\leftarrow\mathbf{1}\{y^{\prime}\neq y\}$. Similar to the classification case, (6) can be estimated through histogram binning. The plot for the function $f_{M}$ is called a reliability diagram (Murphy and Winkler,, 1977) and provides information on overconfidence ($f_{M}(s)<s$) or underconfidence. Importantly, these methods remain applicable in NLG as we have shown. Moreover, they will automatically focus on the task-specific aspects of uncertainty, since the function $f_{M}$ only depends on model generations through the utility $S$.
@@
We discussed the evaluation of LM calibration based on the uncertainty measure (2). To our knowledge, only Huang et al., (2024) studied a similar problem in the general NLG setting, namely the rank-calibration of general uncertainty measures. As discussed in Huang et al., (2024, §4.2), rank-calibration is still a different property; in particular, (2) will be rank-calibrated whenever $f_{M}$ is monotone, even though it can be arbitrarily overconfident or underconfident and thus miscalibrated.

### v2

@@ Abstract
Applications of large language models often involve the generation of free-form responses, in which case uncertainty quantification becomes challenging. This is due to the need to identify task-specific uncertainties (e.g., about the semantics) which appears difficult to define in general cases. This work addresses these challenges from a perspective of Bayesian decision theory, starting from the assumption that our utility is characterized by a similarity measure that compares a generated response with a hypothetical true response. We discuss how this assumption enables principled quantification of the model’s subjective uncertainty and its calibration. We further derive a measure for epistemic uncertainty, based on a missing data perspective and its characterization as an excess risk. The proposed methods can be applied to black-box language models. We illustrate the methods on question answering and machine translation tasks. Our experiments provide a principled evaluation of task-specific calibration, and demonstrate that epistemic uncertainty offers a promising deferral strategy for efficient data acquisition in in-context learning.
@@ 1 Introduction
We are interested in uncertainty quantification (UQ) for language models (LMs) in free-form natural language generation (NLG): given instruction $I$, the model generates a response $y^{\prime}$ based on its predictive distribution $p_{M}(y\mid I)$, where $y^{\prime}$ could be any natural language passage that fits the instruction. An example of this is in question answering (QA): given a question from the user, the model may provide a brief answer, but it may also follow with supporting facts and explanations, which can vary in form and detail. The user can be satisfied by a wide variety of responses, irrespective of their style or (to some extent) the choice of supporting facts included.
@@
This paper is about the observation that the above challenges can be resolved in a more general setup from a perspective of Bayesian decision theory (Savage,, 1954). We start by assuming that we can compute a similarity-type measure $S(y^{\prime},y;I)$ that characterizes our utility function when the model generates $y^{\prime}$ in respond to $I$ and $y$ is a (true or hypothetical) correct response. This generalizes previous works based on semantic equivalence, which corresponds to defining $S$ using entailment. It can also be applied to traditional structured prediction tasks where a choice of $S$ is readily available (e.g., lexical scores for machine translation, Koehn,, 2009), as well as any task where evaluation can be implemented using LMs that compute $S$ based on few-shot demonstrations or detailed instructions (e.g., Yu et al.,, 2023; Wang et al.,, 2023; Bannur et al.,, 2024).
@@
To quantify the subjective uncertainty defined by the LM, we assume the generation $y$ is always chosen to maximize the expected utility under the LM’s predictive distribution $p_{M}$2.1). This is a common, albeit implicit, assumption in NLG (Bickel and Doksum,, 1997; Bertsch et al.,, 2023). It can be supported by a belief that high-capacity LMs may approximate Bayesian inference (e.g., Xie et al.,, 2021; Akyürek et al.,, 2022; Hahn and Goyal,, 2023), and is also more broadly applicable whenever we view $p_{M}$ as our best available model for $y\mid I$. Subjective uncertainty is then naturally characterized by the Bayes risk, or equivalently the maximum achievable expected utility, where the action space is defined by the candidate generations available. This simple observation allows us to understand that previous methods for “semantic uncertainty” can be adapted to a broader range of scenarios, regardless of whether semantic equivalence is relevant. It also highlights a unique, principled approach for aggregating similarity measures among generations (§2.2).
@@
For subjective uncertainty measures to be useful, it will be ideal if the LM is more calibrated. It has been unclear in previous works how calibration can be evaluated in general NLG tasks, given the distinction between relevant and irrelevant differences for a task and our apparent inability to formally characterise it. The decision-theoretic view provides a natural answer to this question: an LM is deemed to possess a calibrated notion of uncertainty if the expected subjective utility of its action matches the expectation of the actually incurred utility w.r.t. the true data distribution (§2.3). This observation allows us to quantify the calibration of LMs through reliability diagrams (Murphy and Winkler,, 1977) and a generalization of the expected calibration error (Naeini et al.,, 2015).
@@
Bayesian modelling enables the decomposition of predictive uncertainty into epistemic uncertainty and aleatoric uncertainty (Der Kiureghian and Ditlevsen,, 2009). Epistemic uncertainty refers to the reducible proportion of uncertainty and can guide data acquisition (Kendall and Gal,, 2017), or more broadly, indicate when additional information can be most effectively used to improve prediction. The quantification of epistemic uncertainty is of interest in many LM applications, as predictions can often benefit from additional information (Lewis et al.,, 2020; Brown et al.,, 2020; Li et al.,, 2023), which may be costly to obtain (Agarwal et al.,, 2024; Min et al.,, 2024). Yet it appears challenging due to the black-box nature of LMs and the free-form generations involved. As we show in §2.4, the decision-theoretic view leads to a principled measure of task-specific epistemic uncertainty in in-context learning (ICL, Brown et al.,, 2020); key additional ingredients include a missing data perspective to Bayesian modelling (Fong et al.,, 2023) and a connection between epistemic uncertainty and an excess risk (Xu and Raginsky,, 2022). The resulted measure is connected to and generalizes several recent methods, which provides independent justification for its use.
The methodological developments allow us to better understand the “probabilistic uncertainty” in modern LMs, that is, the uncertainty reflected in the variation of the model’s predictive distribution. §4 provides experimental illustrations: We revisit past comparisons (Tian et al.,, 2023; Lin et al.,, 2024) of probabilistic and non-probabilistic UQ for instruction-tuned LMs by conducting a principled evaluation of task-specific calibration (§4.1). In §4.2, we demonstrate that epistemic uncertainty provides a deferral strategy (Madras et al.,, 2018; Dohan et al.,, 2022) that facilitates efficient data acquisition in many-shot ICL (Agarwal et al.,, 2024).
@@
The rest of this paper is structured as follows: §2 presents the proposed methodology, §3 reviews related work, §4 presents the experiments, and §5 provides concluding remarks.
2 Quantifying Uncertainty and Calibration
@@ 2.1 A Utilitarian Setup
The setup of this paper is as follows:
@@
  • our utility can be quantified through a similarity measure $S(y^{\prime},y;I)\in\mathbb{R}$, which can be cheaply evaluated;
  • we have an LM with predictive distribution $p_{M}$, and, in the absence of further evidence, consider $p_{M}$ sufficiently trustworthy so that the ideal generation (i.e., action) $y^{\prime}$ for $I$ should maximize the expected utility,
@@
    where $\mathcal{Y}^{\prime}_{I}\subset\mathcal{Y}$ is the space of candidate generations.
In the above, $p_{M}$ represents a subjective belief about the true data distribution. $\mathcal{Y}_{I}^{\prime}\subset\mathcal{Y}$ can be determined based on computational constraints, e.g., as a separate set of samples from $p_{M}(\cdot\mid I)$.111 When $\mathcal{Y}_{I}^{\prime}$ is stochastic, expectation quantities such as (1) should also be averaged over the randomness in $\mathcal{Y}_{I}^{\prime}$. For brevity, we will omit all such (outmost) expectations. The utility $S$ determines a risk function $r(y^{\prime},y;I)=-S(y^{\prime},y;I)$, which can be scaled and shifted as desired.
As discussed in §1, this setup is very general. This decision-theoretic perspective is taken in the framework of minimum Bayes risk (MBR) decoding (Bickel and Doksum,, 1997), which advocates for generating $y$ based precisely on (1). MBR recovers many modern methods for natural language generation (Bertsch et al.,, 2023), which can be viewed as implicitly adopting the same view. Note that the decision-theoretic setup does not compel us to assume that $p_{M}$ is strictly a Bayesian model (e.g., in the sense of Xie et al.,, 2021): although that will provide a good justification, it is also possible that a non-Bayesian LM better approximates our belief than any available Bayesian model does.
2.2 Task-Specific Measures of Subjective Uncertainty
@@
$R_{B}(I):=\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}r(y^{\prime},y;I).$ (2)
For example, if $\mathcal{Y}=\mathcal{Y}^{\prime}_{I}\subset\mathbb{R}$ and $r(y,y^{\prime};I)=(y-y^{\prime})^{2}$, $R_{B}(I)$ will be equivalent to the predictive variance; if with an abuse of notation we redefine $y^{\prime}$ as a density function and $r(y^{\prime},y;I)\leftarrow\log y^{\prime}(y)$, $R_{B,\mathcal{Y}^{\prime}_{I}}(I)$ will become the entropy. When the action space $\mathcal{Y}^{\prime}_{I}\subsetneq\mathcal{Y}$ is a strict subset, (2) becomes connected to measures of “usable information” discussed in Xu et al., (2019).
It is thus natural to also adopt (2) as a measure of subjective uncertainty in NLG. By definitions, it represents the task-specific uncertainty we are interested in, and we recover measures of semantic uncertainty if we define $r$ based on semantic equivalence. (2) can be implemented through Monte-Carlo estimation.
A common generation strategy is defined by the Gibbs predictor, which corresponds to (1) with $\mathcal{Y}^{\prime}_{I}:=\{y_{I}\}$, where $y_{I}\sim p_{M}(\cdot\mid I)$ is a single random sample. In such cases, the expected uncertainty $\mathbb{E}_{\mathcal{Y}^{\prime}_{I}}R_{B}(I)$ is connected to the degree statistic in Lin et al., (2024, Eq. 8); see also Wang et al., (2022, p.8). The difference is that we use a similarity measure that we assume to define our utility. Compared with previous works (Kuhn et al.,, 2023; Lin et al.,, 2024; Nikitin et al.,, 2024) which explored many uncertainty measures based on the aggregation of similarity measures, the decision-theoretic perspective highlights one unique method without any underspecified design choices, justifies its use beyond the scope of semantic equivalence, and shows that different choices of predictors ($\mathcal{Y}_{I}^{\prime}$) should be matched with different uncertainty measures. Although being a measure of subjective uncertainty, it is not guaranteed to always outperform heuristic alternatives on downstream tasks, in particular if the LM is severely miscalibrated.
@@ 2.3 Evaluation of Task-Specific Calibration
Subjective uncertainty measures are most useful when the LM is reasonably calibrated. As discussed in Huang et al., 2024a , many previous works studying uncertainty in free-form NLG did not distinguish between calibration and predictive performance in their evaluation, which leads to a relative lack of understanding about the calibration in such settings.
A common definition of calibration is as follows (see e.g., Johnson et al.,, 2024): suppose the spaces $\mathcal{Y},\mathcal{I}$ are discrete, and denote the true data distribution as $p_{0}$; a model $p_{M}$ is considered calibrated w.r.t. some grouping function $G(I):\mathcal{I}\to\mathcal{G}$ if we have
@@
$G(I)$ can be defined based on uncertainty measures, or it may incorporate additional information such as input demographics. In NLG, however, (3) can be trivially violated if, e.g., the LM always generates a more verbose response compared with $p_{0}$, which may happen due to instruction tuning (Adlakha et al.,, 2023).
In light of the assumptions in §2.1, it is natural to relax (3) to require that for all $g\in\mathcal{G},y^{\prime}\in\mathcal{Y}^{\prime}_{I},s\in\mathbb{R}$
$\mathbb{E}(p_{M}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$ $=\mathbb{E}(p_{0}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$ (3’)
@@
$s_{y,I}:\mathcal{Y}_{I}^{\prime}\to\mathbb{R},~{}y^{\prime}\mapsto S(y^{\prime},y;I)$
is injective. (3) becomes weaker if $S$ only depends on $y$ and $y^{\prime}$ through a non-invertible function, e.g., a map from a natural language passage to its semantics. In such cases, it is clear that (3) remains sufficiently strong while focusing on task-specific calibration.
@@
We can generally view the function $s_{y,I}(y^{\prime})$ as a representation of the “task-specific semantics”.222 For the purposes of MBR generation and uncertainty quantification, it suffices to define the function $s_{y,I}$ on $\mathcal{Y}^{\prime}_{I}$. For other purposes it may be helpful to extend its domain. In fact, the replacement of $y^{\prime}\in\mathcal{Y}^{\prime}_{I}$ with $s_{y^{\prime},I}(\cdot)$ has been advocated by Savage, (1954), and $s_{y,I}$ is referred to as a Savage act (see e.g., Marinacci,, 2015, §2.4).
A main reason we are interested in (3) or (3) is that they guarantee the calibration of uncertainty measures. It follows from (3) that $\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(y^{\prime},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(y^{\prime},y;I)$ for all $y^{\prime}$, and thus
$\mathbb{E}_{p_{0}(I\mid G(I)=g)}R_{B}(I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(\hat{y}_{I},y;I),$
where $\hat{y}_{I}$ denotes the MBR generation (1). Plugging in $G(I)=R_{B}(I)$, we find that the calibration criterion (3) holds only if
$\text{ECE}(p_{M}):=\mathbb{E}_{s}|f_{M}(s)-s|=0,~{}~{}~{}\text{where }f_{M}(s):=\mathbb{E}_{I,y,\hat{y}_{I}}(r(\hat{y}_{I},y;I)\mid R_{B}(I)=s)~{}\forall s\in\mathbb{R}.$ (4)
(4) generalizes the expected calibration error (ECE, Naeini et al.,, 2015) in classification, which it recovers with $r(y^{\prime},y;I)\leftarrow\mathbf{1}\{y^{\prime}\neq y\}$. As in the classification case, we can estimate (4) through histogram binning. The plot for $f_{M}$ is called a reliability diagram (Murphy and Winkler,, 1977) and provides insights on overconfidence ($f_{M}(s)<s$) or underconfidence. Importantly, these methods remain applicable in NLG without the need to introduce any binary notion of correctness. Moreover, they will automatically focus on the task-specific aspects of generation, as ECE and $f_{M}$ only depend on model samples through the utility.